RyleeWarren0000000:

The expression 2(l + w) is used to calculate the perimeter of a rectangle, where l is length and w is width. If the length is Fraction 2 over 3 unit and the width is Fraction 1 over 3 unit, what is the perimeter of the rectangle in units? (3 points) Fraction 2 over 3 unit 1 unit 1 Fraction 2 over 3 units 2 units

JamesTDG:

Your equation is (2/3+1/3)2, 2/3 + 1/3 = 1, so what is 1*2?

RyleeWarren0000000:

your no help man i need the answer

JamesTDG:

I drew it out for you, it is quite simple

snowflake0531:

\[2(l+w)\\2(\frac{2}{3}+\frac{1}{3})\\\frac{2}{3}+\frac{1}{3}=1\\2 \times 1\] so just find out what 2 times 1 is

RyleeWarren0000000:

Evaluate the expression 3.14(a2 + ab) when a = 4 and b = 2. (Input decimals only, such as 12.71, as the answer.) (4 points) whats the answer man